### v1

# Increasing Batch Size Improves Convergence of Stochastic Gradient Descent with Momentum

Abstract
Stochastic gradient descent with momentum (SGDM), which is defined by adding a momentum term to SGD, has been well studied in both theory and practice. Theoretically investigated results showed that the settings of the learning rate and momentum weight affect the convergence of SGDM. Meanwhile, practical results showed that the setting of batch size strongly depends on the performance of SGDM. In this paper, we focus on mini-batch SGDM with constant learning rate and constant momentum weight, which is frequently used to train deep neural networks in practice. The contribution of this paper is showing theoretically that using a constant batch size does not always minimize the expectation of the full gradient norm of the empirical loss in training a deep neural network, whereas using an increasing batch size definitely minimizes it, that is, increasing batch size improves convergence of mini-batch SGDM. We also provide numerical results supporting our analyses, indicating specifically that mini-batch SGDM with an increasing batch size converges to stationary points faster than with a constant batch size. Python implementations of the optimizers used in the numerical experiments are available at https://anonymous.4open.science/r/momentum-increasing-batch-size- [^8] C/.
Machine Learning, ICML
1 Introduction
Stochastic gradient descent (SGD) and its variants, such as SGD with momentum (SGDM) and adaptive methods, are well known as useful optimizers for minimizing the empirical loss defined by the mean of nonconvex loss functions in training a deep neural network (DNN). In the present paper, we focus on SGDM optimizers, which are defined by adding a momentum term to SGD. Various types of SGDM have been proposed, such as stochastic heavy ball (SHB) (Polyak, 1964), normalized-SHB (NSHB) (Gupal & Bazhenov, 1972), Nesterov’s accelerated gradient method (Nesterov, 1983; Sutskever et al., 2013), synthesized Nesterov variants (Lessard et al., 2016), Triple Momentum (Van Scoy et al., 2018), Robust Momentum (Cyrus et al., 2018), PID control-based methods (An et al., 2018), stochastic unified momentum (SUM) (Yan et al., 2018), accelerated SGD (Jain et al., 2018; Kidambi et al., 2018; Varre & Flammarion, 2022; Li et al., 2024), quasi-hyperbolic momentum (QHM) (Ma & Yarats, 2019), and proximal-type SHB (PSHB) (Mai & Johansson, 2020).
Since the empirical loss is nonconvex with respect to a parameter $\bm{\theta}\in\mathbb{R}^{d}$ of a DNN, we are interested in nonconvex optimization for SGDM. Let $\bm{\theta}_{t}\in\mathbb{R}^{d}$ be the $t$-th approximation of SGDM to minimize the nonconvex empirical loss function $f\colon\mathbb{R}^{d}\to\mathbb{R}$. SGDM is defined as $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta_{t}\bm{m}_{t}$, where $\eta_{t}>0$ is a learning rate and $\bm{m}_{t}$ is a momentum buffer. For example, SGDM with $\bm{m}_{t}:=\beta\bm{m}_{t-1}+\nabla f_{B_{t}}(\bm{\theta}_{t})$ is SHB, where $\nabla f_{B_{t}}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ denotes the stochastic gradient of $f$ and $\beta\in[0,1)$ is a momentum weight. SGDM with $\bm{m}_{t}:=\beta\bm{m}_{t-1}+(1-\beta)\nabla f_{B_{t}}(\bm{\theta}_{t})$ is NSHB. Since SHB with $\beta=0$ (NSHB with $\beta=0$) coincides with SGD defined by $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta_{t}\nabla f_{B_{t}}(\bm{\theta}_{t})$, SGDM is defined by adding a momentum term (e.g., $\beta\bm{m}_{t-1}$ in the case of SHB) to SGD.
[…]
Optimizer Gradient Additional Assumption Learning Rate $\eta_{t}$ Weight $\beta_{t}$ Convergence Analysis
(1) SUM Noise Bounded Gradient $\eta=O(\frac{1}{\sqrt{T}})$ $\beta_{t}=\beta$ $\mathbb{E}\|\nabla f_{T}\|=O(\frac{1}{T^{1/4}})$
(2) QHM Noise Bounded Gradient $\eta_{t}\to 0$ $\beta_{t}\to 0$ $\exists(\bm{\theta}_{t_{i}}):\nabla f(\bm{\theta}_{t_{i}})\to 0$
(3) QHM Noise Bounded Gradient $\eta_{t}\to 0$ $\beta_{t}\to 1$ $\exists(\bm{\theta}_{t_{i}}):\nabla f(\bm{\theta}_{t_{i}})\to 0$
(4) PSHB Noise Bounded Gradient $\eta=O(\frac{1}{\sqrt{T}})$ $\beta_{t}=\beta$ $\mathbb{E}\|\nabla f_{T}\|=O(\frac{1}{T^{1/4}})$
(5) SHB Noise ——— $\eta=O(\frac{1}{\sqrt{T}})$ $\beta_{t}=\beta$ $\mathbb{E}\|\nabla f_{T}\|=O(\frac{1}{T^{1/4}})$
(6) NSHB Noise ——— $\eta=O(\frac{1}{L})$ $\beta_{t}=\beta$ $\mathbb{E}\|\nabla f_{T}\|=O(\sqrt{\frac{1}{T}+\sigma^{2}})$
(7) SUM Noise Polyak-Łojasiewicz $\eta_{t}\to 0$ $\beta_{t}=\beta$ $\mathbb{E}[f(\bm{\theta}_{t})]\to f^{\star}$
NSHB Constant ——— $\eta=O(\frac{1}{L})$ $\beta_{t}=\beta$ $\mathbb{E}\|\nabla f_{T}\|=O(\sqrt{\frac{1}{T}+\frac{\sigma^{2}}{b}})$
[Theorem 3.1] Mini-batch
SHB Constant ——— $\eta=O(\frac{1}{L})$ $\beta_{t}=\beta$ $\mathbb{E}\|\nabla f_{T}\|=O(\sqrt{\frac{1}{T}+\frac{\sigma^{2}}{b}})$
[Theorem 3.2] Mini-batch
NSHB Increasing ——— $\eta=O(\frac{1}{L})$ $\beta_{t}=\beta$ $\mathbb{E}\|\nabla f_{T}\|=O(\frac{1}{T^{1/2}})$
[Theorem 3.3] Mini-batch
SHB Increasing ——— $\eta=O(\frac{1}{L})$ $\beta_{t}=\beta$ $\mathbb{E}\|\nabla f_{T}\|=O(\frac{1}{T^{1/2}})$
[Theorem 3.4] Mini-batch
Table 1 summarizes convergence analyses of SGDM for nonconvex optimization. For example, NSHB ((6) in Table 1) using a constant learning rate $\eta_{t}=\eta>0$ and a constant momentum weight $\beta_{t}=\beta$ satisfies $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\sqrt{\frac{1}{T}+\sigma^{2}})$ (Liu et al., 2020, Theorem 1), where $T$ is the number of steps, $\nabla f\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the gradient of $f$, $\sigma^{2}$ is the upper bound of the variance of the stochastic gradient of $f$, and $\mathbb{E}[X]$ denotes the expectation of a random variable $X$. In comparison, QHM ((2) in Table 1), which is a generalization of NSHB, using a decaying learning rate $\eta_{t}$ and a decaying momentum weight $\beta_{t}$ satisfies $\liminf_{t\to+\infty}\|\nabla f(\bm{\theta}_{t})\|=0$ (Gitman et al., 2019, Theorem 1). As can be seen from these convergence analysis results, the performance of SGDM in finding a stationary point $\bm{\theta}^{\star}$ of $f$ (i.e., $\nabla f(\bm{\theta}^{\star})=\bm{0}$) depends on the settings of the learning rate $\eta_{t}$ and the momentum weight $\beta_{t}$.
Moreover, we would like to emphasize that the setting of the batch size $b_{t}$ affects the performance of SGDM. Previous results in (Shallue et al., 2019; Zhang et al., 2019) numerically showed that, for deep learning optimizers, the number of steps needed to train a DNN halves for each doubling of the batch size. In (Smith et al., 2018), it was numerically shown that using an enormous batch size leads to a reduction in the number of parameter updates and the model training time. Hence, we decided to investigate theoretically how the setting of batch size affects convergence of SGDM.
1.1 Contribution
In this paper, we focus on mini-batch SGDM with constant learning rate $\eta>0$ and constant momentum weight $\beta\in[0,1)$, which is frequently used to train DNNs in practice.
  1. 1.
    The first theoretical contribution of the paper is to show that an upper bound of $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]$ for mini-batch SGDM using a constant batch size $b$ is
    $O\left(\sqrt{\frac{f(\bm{\theta}_{0})-f^{\star}}{\eta T}+\frac{L\eta\sigma^{2}}{b}}\right),$
    which implies that mini-batch SGDM does not always minimize the expectation of the full gradient norm of the empirical loss in training a DNN (Table 1; Theorems 3.1 and 3.2).
The bias term $\frac{f(\bm{\theta}_{0})-f^{\star}}{\eta T}$ converges to 00 when $T\to+\infty$. However, the variance term $\frac{L\eta\sigma^{2}}{b}$ remains a constant positive real number regardless of how large $T$ is. In contrast, using a large batch size $b$ makes the variance term $\frac{L\eta\sigma^{2}}{b}$ small. Hence, we can expect that an upper bound of $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]$ for mini-batch SGDM with increasing batch size converges to 00.
  1. 2.
    The second theoretical contribution is to show that an upper bound of $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]$ for mini-batch SGDM with increasing batch size $b_{t}$ such that $b_{t}$ is multiplied by $\delta>1$ every $E$ epochs is
    $O\left(\sqrt{\frac{f(\bm{\theta}_{0})-f^{\star}}{\eta T}+\frac{L\eta\sigma^{2}\delta}{(\beta^{2}\delta-1)b_{0}T}}\right),$ (1)
    which implies that mini-batch SGDM minimizes the expectation of the full gradient norm of the empirical loss in the sense of an $O(\frac{1}{\sqrt{T}})$ rate of convergence (Table 1; Theorems 3.3 and 3.4).
The previous results reported in (Byrd et al., 2012; Balles et al., 2016; De et al., 2017; Smith et al., 2018; Goyal et al., 2018; Shallue et al., 2019; Zhang et al., 2019) indicated that increasing batch sizes is useful for training DNNs with deep learning optimizers. However, providing the theoretical performance of mini-batch SGDM with an increasing batch size may be insufficient, as seen in the existing analyses of SGDM (Table 1). The paper shows theoretically that SGDM with an increasing batch size converges to stationary points of the empirical loss (Theorems 3.3 and 3.4). The previous results in (Yan et al., 2018, Theorem 1), (Mai & Johansson, 2020, Theorem 1), and (Yu et al., 2019, Corollary 1) (Table 1(1), (4), and (5)) showed that SGDM with constant learning rate $\eta=O(\frac{1}{\sqrt{T}})$ and a constant momentum weight $\beta$ has convergence rate $O(\frac{1}{T^{1/4}})$. Our results (Theorems 3.3 and 3.4) guarantee that, if the batch size increases, then SGDM satisfies $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\frac{1}{T^{1/2}})$, which improves the previous convergence rate $O(\frac{1}{T^{1/4}})$.
The result in (1) indicates that the performance of mini-batch SGDM with increasing batch size $b_{t}$ depends on $\delta$. Let $\eta$ and $\beta$ be fixed (e.g., $\eta=0.1$ and $\beta=0.9$). Then, (1) indicates that the larger $\delta$ is, the smaller the variance term $\frac{L\eta\sigma^{2}\delta}{(\beta^{2}\delta-1)b_{0}T}$ is (since $\frac{\delta}{\beta^{2}\delta-1}=\frac{1}{(0.9)^{2}-1/\delta}$ becomes small as $\delta$ becomes large). We are interested in verifying whether this theoretical result holds in practice.
  1. 3.
    The third contribution is showing numerically that quadruply increasing batch size (i.e., $\delta=4$) decreases $\min_{t\in[0:T-1]}\|\nabla f(\bm{\theta}_{t})\|$ faster than doubly increasing batch size (i.e., $\delta=2$) or maintaining a constant batch size.
We consider training ResNet-18 on the CIFAR-100 and Tiny ImageNet datasets using not only NSHB and SHB but also baseline optimizers: SGD, Adam (Kingma & Ba, 2015), AdamW (Loshchilov & Hutter, 2019), and RMSprop (Tieleman & Hinton, 2012). A particularly interesting result in Section 4 is that an increasing batch size is applicable for Adam in the sense of minimizing $\min_{t\in[0:T-1]}\|\nabla f(\bm{\theta}_{t})\|$ fastest. Hence, in the future, we would like to verify whether Adam with an increasing batch size theoretically has a better convergence rate than SGDM.
2 Mini-batch SGDM for Empirical Risk Minimization
2.1 Empirical risk minimization
[…]
We assume that the loss functions $f_{i}$ ($i\in[n]$) satisfy the conditions in the following assumption.
Assumption 2.1.
Let $n$ be the number of training samples and let $L_{i}>0$ ($i\in[n]$).
(A1) $f_{i}\colon\mathbb{R}^{d}\to\mathbb{R}$ ($i\in[n]$) is differentiable and $L_{i}$-smooth (i.e., there exists $L_{i}>0$ such that, for all $\bm{\theta}_{1},\bm{\theta}_{2}\in\mathbb{R}^{d}$, $\|\nabla f_{i}(\bm{\theta}_{1})-\nabla f_{i}(\bm{\theta}_{2})\|\leq L_{i}\|\bm{\theta}_{1}-\bm{\theta}_{2}\|$), $L:=\frac{1}{n}\sum_{i\in[n]}L_{i}$, and $f^{\star}$ is the minimum value of $f$ over $\mathbb{R}^{d}$.
(A2) Let $\xi$ be a random variable independent of $\bm{\theta}\in\mathbb{R}^{d}$. $\nabla f_{\xi}\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the stochastic gradient of $\nabla f$ such that (i) for all $\bm{\theta}\in\mathbb{R}^{d}$, $\mathbb{E}_{\xi}[\nabla f_{\xi}(\bm{\theta})]=\nabla f(\bm{\theta})$ and (ii) there exists $\sigma\geq 0$ such that, for all $\bm{\theta}\in\mathbb{R}^{d}$, $\mathbb{V}_{\xi}[\nabla f_{\xi}(\bm{\theta})]=\mathbb{E}_{\xi}[\|\nabla f_{\xi}(\bm{\theta})-\nabla f(\bm{\theta})\|^{2}]\leq\sigma^{2}$, where $\mathbb{E}_{\xi}[\cdot]$ denotes expectation with respect to $\xi$.
(A3) Let $b\in\mathbb{N}$ such that $b\leq n$ and let $\bm{\xi}=(\xi_{1},\xi_{2},\cdots,\xi_{b})^{\top}$ comprise $b$ independent and identically distributed variables and be independent of $\bm{\theta}\in\mathbb{R}^{d}$. The full gradient $\nabla f(\bm{\theta})$ is estimated as the mini-batch gradient at $\bm{\theta}$ defined by $\nabla f_{B}(\bm{\theta}):=\frac{1}{b}\sum_{i=1}^{b}\nabla f_{\xi_{i}}(\bm{\theta})$.
2.2 Mini-batch NSHB and mini-batch SHB
Let $\bm{\theta}_{t}\in\mathbb{R}^{d}$ be the $t$-th approximated parameter of DNN. Then, mini-batch NSHB uses $b_{t}$ loss functions $f_{\xi_{t,1}},\cdots,f_{\xi_{t,b_{t}}}$ randomly chosen from $\{f_{1},\cdots,f_{n}\}$ at each step $t$, where $\bm{\xi}_{t}=(\xi_{t,1},\cdots,\xi_{t,b_{t}})^{\top}$ is independent of $\bm{\theta}_{t}$ and $b_{t}$ is a batch size satisfying $b_{t}\leq n$. The Mini-batch NSHB optimizer is as in Algorithm 1.
Algorithm 1 Mini-batch NSHB optimizer
0:  $\bm{\theta}_{0},\bm{m}_{-1}:=\bm{0}$ (initial point), $b_{t}>0$ (batch size), $\eta>0$ (learning rate), $\beta\in[0,1)$ (momentum weight), $T\geq 1$ (steps)
0:  $(\bm{\theta}_{t})\subset\mathbb{R}^{d}$
1:  for $t=0,1,\ldots,T-1$ do
2:      $\nabla f_{B_{t}}(\bm{\theta}_{t}):=\frac{1}{b_{t}}\sum_{i=1}^{b_{t}}\nabla f_{\xi_{t,i}}(\bm{\theta}_{t})$
3:      $\bm{m}_{t}:=\beta\bm{m}_{t-1}+(1-\beta)\nabla f_{B_{t}}(\bm{\theta}_{t})$
4:      $\bm{\theta}_{t+1}:=\bm{\theta}_{t}-\eta\bm{m}_{t}$
5:  end for
The simplest optimizer for adding a momentum term (denoted by $\beta\bm{m}_{t-1}$) to SGD is the stochastic heavy ball (SHB) method (Polyak, 1964), which is provided in PyTorch (Paszke et al., 2019). SHB is defined as follows:
$\begin{split}\bm{m}_{t}=\beta\bm{m}_{t-1}+\nabla f_{B_{t}}(\bm{\theta}_{t}),\text{ }\bm{\theta}_{t+1}=\bm{\theta}_{t}-\alpha\bm{m}_{t},\end{split}$ (2)
where $\beta\in[0,1)$ and $\alpha>0$. SHB defined by (2) with $\beta=0$ coincides with SGD. SHB defined by (2) has the form $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\alpha\nabla f_{B_{t}}(\bm{\theta}_{t})+\beta(\bm{\theta}_{t}-\bm{\theta}_{t-1})$. Meanwhile, Algorithm 1 is called the normalized-SHB (NSHB) optimizer (Gupal & Bazhenov, 1972) and has the form $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta(1-\beta)\nabla f_{B_{t}}(\bm{\theta}_{t})+\beta(\bm{\theta}_{t}-\bm{\theta}_{t-1})$. Hence, NSHB (Algorithm 1) with $\eta=\frac{\alpha}{1-\beta}$ coincides with SHB defined by (2).
3 Mini-batch SGDM with Constant and Increasing Batch Sizes
3.1 Constant batch size scheduler
The following indicates that an upper bound of $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]$ of mini-batch NSHB using a constant batch size
[…]
does not always converge to 00 (a proof of Theorem 3.1 is given in Appendix A.3).
Theorem 3.1 (Upper bound of $\min_{t}\mathbb{E}\|\nabla f(\bm{\theta}_{t})\|$ of mini-batch NSHB with Constant BS).
Suppose that Assumption 2.1 holds and consider the sequence $(\bm{\theta}_{t})$ generated by Algorithm 1 with a momentum weight $\beta\in(0,1)$, a constant learning rate $\eta>0$ such that
$\eta\leq\max\left\{\frac{1-\beta}{2\sqrt{2}\sqrt{\beta+\beta^{2}}L},\frac{(1-\beta)^{2}}{(5-\beta+2\beta^{2})L}\right\},$
and Constant BS defined by (3), where $L:=\frac{1}{n}\sum_{i\in[n]}L_{i}$ and $f^{\star}$ is the minimum value of $f$ over $\mathbb{R}^{d}$ (see (A1)). Then, for all $T\geq 1$,
$\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
$\leq\frac{2(f(\bm{\theta}_{0})-f^{\star})}{\eta T}+\frac{L\eta\sigma^{2}}{b}\left\{\frac{3\beta^{2}+\beta}{2(1+\beta)}+1\right\},$
that is,
$\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O\left(\sqrt{\frac{1}{T}+\frac{\sigma^{2}}{b}}\right).$
From the discussion in Section 2.2 such that NSHB (Algorithm 1) with $\eta=\frac{\alpha}{1-\beta}$ coincides with SHB defined by (2), Theorem 3.1 leads to the following.
Theorem 3.2 (Upper bound of $\min_{t}\mathbb{E}\|\nabla f(\bm{\theta}_{t})\|$ of mini-batch SHB with Constant BS).
Suppose that Assumption 2.1 holds and consider the sequence $(\bm{\theta}_{t})$ generated by (2) with a momentum weight $\beta\in(0,1)$, a constant learning rate $\alpha>0$ such that
$\alpha\leq\max\left\{\frac{(1-\beta)^{2}}{2\sqrt{2}\sqrt{\beta+\beta^{2}}L},\frac{(1-\beta)^{3}}{(5-\beta+2\beta^{2})L}\right\},$
and Constant BS defined by (3). Then, for all $T\geq 1$,
$\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
$\leq\frac{2(1-\beta)(f(\bm{\theta}_{0})-f^{\star})}{\alpha T}+\frac{L\alpha\sigma^{2}}{(1-\beta)b}\left\{\frac{3\beta^{2}+\beta}{2(1+\beta)}+1\right\},$
that is,
$\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O\left(\sqrt{\frac{1}{T}+\frac{\sigma^{2}}{b}}\right).$
3.2 Increasing batch size scheduler
We consider an increasing batch size $b_{t}$ such that
$b_{t}\leq b_{t+1}\text{ }(t\in\mathbb{N}).$
An example of $b_{t}$ (Smith et al., 2018; Umeda & Iiduka, 2024) is, for all $m\in[0:M]$ and all $t\in S_{m}=\mathbb{N}\cap[\sum_{k=0}^{m-1}K_{k}E_{k},\sum_{k=0}^{m}K_{k}E_{k})$ ($S_{0}:=\mathbb{N}\cap[0,K_{0}E_{0})$),
[…]
where $\delta>1$, and $E_{m}$ and $K_{m}$ are the numbers of, respectively, epochs and steps per epoch when the batch size is $\delta^{m}b_{0}$. For example, the exponential growth batch size defined by (4) with $\delta=2$ makes batch size double each $E_{m}$ epochs. We may modify the parameters $a$ and $\delta$ to $a_{t}$ and $\delta_{t}$ monotone increasing with $t$. The total number of steps for the batch size to increase $M$ times is $T=\sum_{m=0}^{M}K_{m}E_{m}$.
The following is a convergence analysis of Algorithm 1 with increasing batch sizes.
Theorem 3.3 (Convergence of mini-batch NSHB with Exponential Growth BS).
Suppose that Assumption 2.1 holds and consider the sequence $(\bm{\theta}_{t})$ generated by Algorithm 1 with a momentum weight $\beta\in(0,1)$, a constant learning rate $\eta>0$ such that
$\eta\leq\max\left\{\frac{1-\beta}{2\sqrt{2}\sqrt{\beta+\beta^{2}}L},\frac{(1-\beta)^{2}}{(5\beta^{2}-6\beta+5)L}\right\},$ (5)
and Exponential Growth BS defined by (4) with $\delta>1$ and $\beta^{2}\delta>1$. Then, for all $T\geq 1$,
$\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\leq\frac{2(f(\bm{\theta}_{0})-f^{\star})}{\eta T}$
$+\frac{2L\eta\sigma^{2}K_{\max}E_{\max}\delta}{(\beta^{2}\delta-1)b_{0}T}\left(\frac{\beta^{2}}{1-\beta^{2}}-\frac{1}{\delta-1}\right),$
where $K_{\max}:=\max\{K_{m}\colon m\in[0:M]\}$ and $E_{\max}:=\max\{E_{m}\colon m\in[0:M]\}$, that is,
$\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O\left(\frac{1}{\sqrt{T}}\right).$
From the discussion in Section 2.2 such that NSHB (Algorithm 1) with $\eta=\frac{\alpha}{1-\beta}$ coincides with SHB defined by (2), Theorem 3.3 leads to the following convergence rate of SHB defined by (2) with an increasing batch size.
Theorem 3.4 (Convergence of mini-batch SHB with Exponential Growth BS).
Suppose that Assumption 2.1 holds and consider the sequence $(\bm{\theta}_{t})$ generated by (2) with a momentum weight $\beta\in(0,1)$, a constant learning rate $\alpha>0$ such that
$\alpha\leq\max\left\{\frac{(1-\beta)^{2}}{2\sqrt{2}\sqrt{\beta+\beta^{2}}L},\frac{(1-\beta)^{3}}{(5\beta^{2}-6\beta+5)L}\right\},$
and Exponential Growth BS defined by (4) with $\delta>1$ and $\beta^{2}\delta>1$. Then, for all $T\geq 1$,
$\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\leq\frac{2(1-\beta)(f(\bm{\theta}_{0})-f^{\star})}{\alpha T}$
$+\frac{2L\alpha\sigma^{2}K_{\max}E_{\max}\delta}{(1-\beta)(\beta^{2}\delta-1)b_{0}T}\left(\frac{\beta^{2}}{1-\beta^{2}}-\frac{1}{\delta-1}\right),$
that is,
$\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O\left(\frac{1}{\sqrt{T}}\right).$
Here, we sketch a proof of Theorem 3.3 (a detailed proof of Theorem 3.3 is given in Appendix A.2).
  1. 1.
    We first show that an upper bound of the variance of $\nabla f_{B_{t}}(\bm{\theta}_{t})$ is $\frac{\sigma^{2}}{b_{t}}$ (Proposition A.1) and then that an upper bound of the variance of $\bm{m}_{t}=(1-\beta)\sum_{i=0}^{t}\beta^{t-i}\nabla f_{B_{i}}(\bm{\theta}_{i})$ is $(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}$ (Lemma A.2) using the idea of the proof of (Liu et al., 2020, Lemma 1).
  2. 2.
    We next show that an auxiliary point $\bm{z}_{t}=\frac{1}{1-\beta}\bm{\theta}_{t}-\frac{\beta}{1-\beta}\bm{\theta}_{t-1}$ ($t\geq 1$), which is used to analyze SGDM (Yan et al., 2018; Yu et al., 2019; Liu et al., 2020), satisfies $\mathbb{E}_{\bm{\xi}_{t}}[f(\boldsymbol{z}_{t+1})]\leq f(\boldsymbol{z}_{t})-\eta\underbrace{\mathbb{E}_{\bm{\xi}_{t}}[\langle\nabla f(\boldsymbol{z}_{t}),\nabla f_{B_{t}}(\bm{\theta}_{t})\rangle]}_{X_{t}}+\frac{L\eta^{2}}{2}\underbrace{\mathbb{E}_{\bm{\xi}_{t}}[\|\nabla f_{B_{t}}(\bm{\theta}_{t})\|^{2}]}_{Y_{t}}$ using the descent lemma (see (7)). Using the Cauchy–Schwarz inequality, Young’s inequality, and the upper bound of the variance of $\bm{m}_{t}$ (Lemma A.2) provides an upper bound of $-\eta\mathbb{E}[X_{t}]$ (see (12)). An upper bound of $\mathbb{E}[Y_{t}]$ is provided by using the upper bound $\frac{\sigma^{2}}{b_{t}}$ of the variance of $\nabla f_{B_{t}}(\bm{\theta}_{t})$ (see Lemma A.3 for details of the upper bounds of $-\eta\mathbb{E}[X_{t}]$ and $\mathbb{E}[Y_{t}]$).
  3. 3.
    We define the Lyapunov function $L_{t}$ by $L_{t}=f(\bm{z}_{t})-f^{\star}+\sum_{i=1}^{t-1}c_{i}\|\bm{\theta}_{t+1-i}-\bm{\theta}_{t-i}\|^{2}$, where $c_{i}$ is defined as in Lemma A.4. Using the above upper bounds of $-\eta\mathbb{E}[X_{t}]$ and $\mathbb{E}[Y_{t}]$, we have that $\mathbb{E}[L_{t+1}-L_{t}]\leq-D\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]+U_{t}$ (Lemma A.4), where $D\in\mathbb{R}$ depends on $\eta$, $\beta$, and $c_{1}$, and $U_{t}>0$ depends on $\sigma^{2}$, $b_{t}$, and $c_{1}$.
  4. 4.
    Setting $\eta$ to satisfy (5) leads to the finding that $D\geq\frac{\eta}{2}>0$ and $U_{t}\leq L\eta^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}$ (see (19) and (20)). As a result, we have that
    $\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$ $\leq\frac{2L_{0}}{\eta T}+\frac{2L\eta\sigma^{2}}{T}\sum_{t=0}^{T-1}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}$
    (see Lemma A.5). Finally, using (4) leads to the assertion of Theorem 3.3.
3.3 Setting of hyperparameter $\delta$ in Exponential Growth BS (4)
Let $\eta$ and $\beta$ be fixed in Algorithm 1 (e.g., $\eta=0.1$ and $\beta=0.9$). Then, Theorems 3.3 and 3.4 indicate that an upper bound of $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]$ for each of mini-batch NSHB and mini-batch SHB with Exponential Growth BS (4) is $O(\sqrt{\frac{f(\bm{\theta}_{0})-f^{\star}}{\eta T}+\frac{L\eta\sigma^{2}\delta}{(\beta^{2}\delta-1)b_{0}T}})$, which implies that the larger $\delta$ is, the smaller the variance term $\frac{L\eta\sigma^{2}\delta}{(\beta^{2}\delta-1)b_{0}T}$ is (since $\frac{\delta}{\beta^{2}\delta-1}=\frac{1}{(0.9)^{2}-1/\delta}$ becomes small as $\delta$ becomes large). In Section 4, we verify whether this theoretical result holds in practice.
3.4 Comparisons of our convergence results with previous ones
Let us compare Theorems 3.13.4 with the previous results listed in Table 1. Theorem 1 in (Liu et al., 2020) ((6) in Table 1) indicated that NSHB using a constant learning rate $\eta=O(\frac{1}{L})$ and a constant momentum weight $\beta$ satisfies $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\sqrt{\frac{1}{T}+\sigma^{2}})$. Since the upper bound $O(\sqrt{\frac{1}{T}+\sigma^{2}})$ converges to $O(\sigma)>0$ when $T\to+\infty$, NSHB in this case does not always converge to stationary points of $f$. The result in (Liu et al., 2020) coincides with Theorem 3.1 indicating that NSHB has $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\sqrt{\frac{1}{T}+\frac{\sigma^{2}}{b}})$ in the sense that NSHB using a constant learning rate and momentum weight does not converge to stationary points of $f$111(Liu et al., 2020, Theorem 3) was a convergence of multistage SGDM. However, since the proof of (Liu et al., 2020, (60), Pages 35 and 36) might not hold for $\beta_{i}<1$, the theorem does not apply here.. Corollary 1 in (Yu et al., 2019) ((5) in Table 1) indicated that SHB using constant $\eta=O(\frac{1}{\sqrt{T}})$ and constant momentum weight $\beta$ satisfies $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\frac{1}{T^{1/4}})$. Using $\eta=O(\frac{1}{\sqrt{T}})$ is necessary to set the number of steps $T$ before implementing SHB. Since the $T$ is fixed, we cannot diverge $T$, that is, the upper bound $O(\frac{1}{T^{1/4}})$ for SHB is a fixed positive constant and does not converge to 00. Meanwhile, Theorem 3.2 is the result for SHB using a constant learning rate $\eta=O(\frac{1}{L})$ and a constant momentum weight, which can be obtained by Theorem 3.1. Theorem 3.2 indicates that SHB has $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\sqrt{\frac{1}{T}+\frac{\sigma^{2}}{b}})$. Hence, Theorem 3.2 coincides with the result in Corollary 1 in (Yu et al., 2019) in the sense that SHB with constant learning rate does not always converge to stationary points of $f$.
Theorems 1 and 2 in (Gitman et al., 2019) ((2) and (3) in Table 1) indicated that QHM, which is a generalization of NSHB, using a decaying learning rate $\eta_{t}$ and a decaying momentum weight $\beta_{t}$ or an increasing momentum weight $\beta_{t}$ satisfies $\liminf_{t\to+\infty}\|\nabla f(\bm{\theta}_{t})\|=0$. Our results in Theorems 3.3 and 3.4 guarantee the convergence of NSHB and SHB with constant learning rate $\eta=O(\frac{1}{L})$, constant momentum weight $\beta$, and an increasing batch size $b_{t}$ in the sense of $\min_{t\in[0:T-1]}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|]=O(\frac{1}{\sqrt{T}})$.
4 Numerical Results
We examined training ResNet-18 on the CIFAR-100 dataset using not only NSHB and SHB but also baseline optimizers: SGD, Adam, AdamW, and RMSprop with constant and increasing batch sizes (see Appendix A.4 for training ResNet-18 on Tiny ImageNet). We used a computer equipped with NVIDIA A100 80GB and Dual Intel Xeon Silver 4316 2.30GHz, 40 Cores (20 cores per CPU, 2 CPUs). The software environment was Python 3.8.2, PyTorch 2.2.2+cu118, and CUDA 12.2. We set the total number of epochs to $E=200$ and the constant momentum weight as the default values in PyTorch. The learning rate was set for Adam and AdamW to $10^{-3}$, for RMSprop to $10^{-2}$, and for SGD, SHB, and NSHB to $10^{-1}$; see also Figure 1(a).
Let us first consider the learning rate and batch size scheduler in Figure 1(a) with a constant batch size ($b=2^{7}$). Figure 1(b) compares the full gradient norm $\min_{e\in[E]}\|\nabla f(\bm{\theta}_{e})\|$ for training for each optimizer and indicates that SHB decreased the full gradient norm quickly. Figures 1(c) and (d) compare the empirical loss $f(\bm{\theta}_{e})$ and the test accuracy score. These figures indicate that SGD, SHB, and NSHB minimized $f$ quickly and had test accuracies of approximately 70 %. Next, let us compare Figure 1 with Figure 2 under the scheduler with the same learning rates in Figure 1(a) and doubly batch size every 20 epochs with the initial batch size $b_{0}=2^{3}$. Figures 2(b) and (c) both show that using a doubly increasing batch size results in a faster decrease of $\min_{e\in[E]}\|\nabla f(\bm{\theta}_{e})\|$ and $f(\bm{\theta}_{e})$, compared to using a constant batch size as in Figures 1(b) and (c). The numerical results in Figures 1(b) and 2(b) are supported theoretically by Theorems 3.13.4 indicating that NSHB and SHB with increasing batch sizes minimize the gradient norm of $f$ faster than with constant batch sizes. In Figures 1(d) and 2(d), it can be seen that using a doubly increasing batch size leads to improved test accuracy for all optimizers except SHB, compared to using a constant batch size. Earlier, we observed that, with a constant batch size, convergence is slower, and accuracy improvement is more gradual. On the other hand, these results suggest that using an increasing batch size leads to a faster convergence and more efficient training. Additionally, when using an increasing batch size, the optimizer’s performance is better overall, particularly in terms of having faster convergence.
Now, let us compare Figure 2 ($\delta=2$) with Figure 3 ($\delta=4$) under the scheduler with the same learning rates as in Figure 1(a) and the batch size quadruply increasing every 40 epochs with an initial batch size $b_{0}=2^{3}$. From Figures 3(b) and (c), it can be observed that the larger the batch size is, the faster the decrease of the full gradient norm $\|\nabla f(\bm{\theta}_{e})\|$ and the empirical loss $f(\bm{\theta}_{e})$ are. Specifically, the quadruply increasing batch size ($\delta=4$; Figure 3) decreases the full gradient norm $\|\nabla f(\bm{\theta}_{e})\|$ and the empirical loss $f(\bm{\theta}_{e})$ more rapidly than the doubly increasing batch size ($\delta=2$; Figure 2). Figures 2(d) and 3(d) indicate that SGD and NSHB had greater than 70 % test accuracies, which implies in turn that, for SGD and NSHB, using an increasing batch size would improve generalization more than using a constant batch size (Figure 1(d)).
4.1 Discussion and future work
Fast convergence of Adam: A particularly interesting result in Figures 23 is that an increasing batch size is applicable for Adam in the sense of minimizing the full gradient norm of $f$ fastest. Hence, we can expect that Adam with an increasing batch size has a convergence rate better than the $O(\frac{1}{\sqrt{T}})$ convergence rate of NSHB and SHB in Theorems 3.3 and 3.4. In the future, we should verify the result holds theoretically.
Full gradient norm and training loss versus test accuracy: As promised in Theorems 3.1 and 3.3, NSHB with increasing batch sizes ($\delta=2,4$) minimized the full gradient norm of $f$ faster than with a constant batch size (Figures 1(b), 2(b), and 3(b)). As a result, NSHB with an increasing batch size ($\delta=2,4$) minimized the training loss $f$ (Figures 1(c), 2(c), and 3(c)) and had higher test accuracies than with a constant batch size (Figures 1(d), 2(d), and 3(d)). Moreover, Figures 13 indicate that AdamW had almost the same trend. Although Adam and AdamW with increasing batch sizes minimized $f$ quickly, the test accuracy of Adam was different from the test accuracy of AdamW (Figure 3(d)). Here, we have the following insights:
  • An increasing batch size quickly minimizes the full gradient norm of the training loss in both theory and practice. In particular, SGDM with an increasing batch size converges to stationary points of the training loss, as promised in our theoretical results.
  • Optimal increasing batch size schedulers with which optimizers have high test accuracies should be discussed. Specifically, we need to discuss optimal $E_{m}$ and $\delta$ such that SGDM and adaptive methods (e.g., Adam and AdamW) improve generalization.
5 Conclusion
This paper presented convergence analyses of mini-batch SGDM with a constant learning rate and momentum weight. Using a constant batch size does not lead to convergence of mini-batch SGDM to stationary points of the training loss, but using an increasing batch size does lead to its convergence. This paper also provided numerical results to support our convergence analyses. In particular, using a quadruply increasing batch size had faster convergence of mini-batch SGDM than using a doubly increasing batch size. Moreover, the numerical results indicated that using an increasing batch size is also applicable for adaptive methods, such as Adam and AdamW, in the sense of minimizing the full gradient norm of the training loss. Hence, in the future, we need to verify theoretically whether adaptive methods with increasing batch sizes have faster convergence than SGDM.
(a) Learning rate and batch size versus epochs
(b) Full gradient norm versus epochs
(c) Empirical loss versus epochs
(d) Test accuracy score versus epochs
Figure 1: (a) Schedulers for each optimizer with constant learning rates and constant batch size, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on CIFAR-100 dataset.
(a) Learning rate and batch size versus epochs
(b) Full gradient norm versus epochs
(c) Empirical loss versus epochs
(d) Test accuracy score versus epochs
Figure 2: (a) Schedulers for each optimizer with constant learning rate and batch size doubly increasing every 20 epochs, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on CIFAR-100 dataset.
(a) Learning rate and batch size versus epochs
(b) Full gradient norm versus epochs
(c) Empirical loss versus epochs
(d) Test accuracy score versus epochs
Figure 3: (a) Schedulers for each optimizer with constant learning rates and batch size quadruply increasing every 40 epochs, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on CIFAR-100 dataset.
Appendix A Proofs of Theorems in the Paper
A.1 Proposition and Lemma
The following proposition holds for the mini-batch gradient.
Proposition A.1.
Let $t\in\mathbb{N}$, $\bm{\xi}_{t}$ be a random variable independent of $\bm{\xi}_{j}$ ($j\in[0:t-1]$), $\bm{\theta}_{t}\in\mathbb{R}^{d}$ be independent of $\bm{\xi}_{t}$, and $\nabla f_{B_{t}}(\bm{\theta}_{t})$ be the mini-batch gradient, where $f_{\xi_{t,i}}$ ($i\in[b_{t}]$) is the stochastic gradient (see Assumption 2.1(A2)). Then, the following hold:
$\mathbb{E}_{\bm{\xi}_{t}}\left[\nabla f_{B_{t}}(\bm{\theta}_{t})\Big{|}\hat{\bm{\xi}}_{t-1}\right]=\nabla f(\bm{\theta}_{t})\text{ and }\mathbb{V}_{\bm{\xi}_{t}}\left[\nabla f_{B_{t}}(\bm{\theta}_{t})\Big{|}\hat{\bm{\xi}}_{t-1}\right]\leq\frac{\sigma^{2}}{b_{t}},$
where $\mathbb{E}_{\bm{\xi}_{t}}[\cdot|\hat{\bm{\xi}}_{t-1}]$ and $\mathbb{V}_{\bm{\xi}_{t}}[\cdot|\hat{\bm{\xi}}_{t-1}]$ are respectively the expectation and variance with respect to $\bm{\xi}_{t}$ conditioned on $\bm{\xi}_{t-1}=\hat{\bm{\xi}}_{t-1}$.
Proof.
Assumption 2.1(A3) and the independence of $b_{t}$ and $\bm{\xi}_{t}$ ensure that
$\mathbb{E}_{\bm{\xi}_{t}}\left[\nabla f_{B_{t}}(\bm{\theta}_{t})\Big{|}\hat{\bm{\xi}}_{t-1}\right]=\mathbb{E}_{\bm{\xi}_{t}}\left[\frac{1}{b_{t}}\sum_{i=1}^{b_{t}}\nabla f_{\xi_{t,i}}(\bm{\theta}_{t})\Bigg{|}\hat{\bm{\xi}}_{t-1}\right]=\frac{1}{b_{t}}\sum_{i=1}^{b_{t}}\mathbb{E}_{\xi_{t,i}}\left[\nabla f_{\xi_{t,i}}(\bm{\theta}_{t})\Big{|}\hat{\bm{\xi}}_{t-1}\right],$
which, together with Assumption 2.1(A2)(i) and the independence of $\bm{\xi}_{t}$ and $\bm{\xi}_{t-1}$, implies that
$\mathbb{E}_{\bm{\xi}_{t}}\left[\nabla f_{B_{t}}(\bm{\theta}_{t})\Big{|}\hat{\bm{\xi}}_{t-1}\right]=\frac{1}{b_{t}}\sum_{i=1}^{b_{t}}\nabla f(\bm{\theta}_{t})=\nabla f(\bm{\theta}_{t}).$ (6)
Assumption 2.1(A3), the independence of $b_{t}$ and $\bm{\xi}_{t}$, and (6) imply that
$\mathbb{V}_{\bm{\xi}_{t}}\left[\nabla f_{B_{t}}(\bm{\theta}_{t})\Big{|}\hat{\bm{\xi}}_{t-1}\right]$ $=\mathbb{E}_{\bm{\xi}_{t}}\left[\|\nabla f_{B_{t}}(\bm{\theta}_{t})-\nabla f(\bm{\theta}_{t})\|^{2}\Big{|}\hat{\bm{\xi}}_{t-1}\right]$
$=\mathbb{E}_{\bm{\xi}_{t}}\left[\left\|\frac{1}{b_{t}}\sum_{i=1}^{b_{t}}\nabla f_{\xi_{t,i}}(\bm{\theta}_{t})-\nabla f(\bm{\theta}_{t})\right\|^{2}\Bigg{|}\hat{\bm{\xi}}_{t-1}\right]$
$=\frac{1}{b_{t}^{2}}\mathbb{E}_{\bm{\xi}_{t}}\left[\left\|\sum_{i=1}^{b_{t}}\left(\nabla f_{\xi_{t,i}}(\bm{\theta}_{t})-\nabla f(\bm{\theta}_{t})\right)\right\|^{2}\Bigg{|}\hat{\bm{\xi}}_{t-1}\right].$
From the independence of $\xi_{t,i}$ and $\xi_{t,j}$ ($i\neq j$) and Assumption 2.1(A2)(i), for all $i,j\in[b_{t}]$ such that $i\neq j$,
$\mathbb{E}_{\xi_{t,i}}[\langle\nabla f_{\xi_{t,i}}(\bm{\theta}_{t})-\nabla f(\bm{\theta}_{t}),\nabla f_{\xi_{t,j}}(\bm{\theta}_{t})-\nabla f(\bm{\theta}_{t})\rangle|\hat{\bm{\xi}}_{t-1}]$
$=\langle\mathbb{E}_{\xi_{t,i}}[\nabla f_{\xi_{t,i}}(\bm{\theta}_{t})|\hat{\bm{\xi}}_{t-1}]-\mathbb{E}_{\xi_{t,i}}[\nabla f(\bm{\theta}_{t})|\hat{\bm{\xi}}_{t-1}],\nabla f_{\xi_{t,j}}(\bm{\theta}_{t})-\nabla f(\bm{\theta}_{t})\rangle$
$=0.$
Hence, Assumption 2.1(A2)(ii) guarantees that
$\mathbb{V}_{\bm{\xi}_{t}}\left[\nabla f_{B_{t}}(\bm{\theta})\Big{|}\hat{\bm{\xi}}_{t-1}\right]=\frac{1}{b_{t}^{2}}\sum_{i=1}^{b_{t}}\mathbb{E}_{\xi_{t,i}}\left[\left\|\nabla f_{\xi_{t,i}}(\bm{\theta}_{t})-\nabla f(\bm{\theta}_{t})\right\|^{2}\Big{|}\hat{\bm{\xi}}_{t-1}\right]\leq\frac{\sigma^{2}b_{t}}{b_{t}^{2}}=\frac{\sigma^{2}}{b_{t}},$
which completes the proof. ∎
Motivated by Lemma 1 in (Liu et al., 2020), we prove the following lemma.
Lemma A.2.
Under Assumption 2.1, Algorithm 1 satisfies that, for all $t\in\{0\}\cup\mathbb{N}$,
$\mathbb{E}\left[\left\|\boldsymbol{m}_{t}-\left(1-\beta\right)\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}\right]\leq(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}},$
where $\mathbb{E}$ denotes the total expectation defined by $\mathbb{E}:=\mathbb{E}_{\bm{\xi}_{0}}\mathbb{E}_{\bm{\xi}_{1}}\cdots\mathbb{E}_{\bm{\xi}_{t}}$.
Proof.
The definition of $\bm{m}_{t}$ and $\bm{m}_{-1}:=\bm{0}$ ensure that
$\bm{m}_{t}$ $=\beta\bm{m}_{t-1}+(1-\beta)\nabla f_{B_{t}}(\bm{\theta}_{t})$
$=\beta\{\beta\bm{m}_{t-2}+(1-\beta)\nabla f_{B_{t-1}}(\bm{\theta}_{t-1})\}+(1-\beta)\nabla f_{B_{t}}(\bm{\theta}_{t})$
$=\beta^{2}\bm{m}_{t-2}+(1-\beta)\{\beta\nabla f_{B_{t-1}}(\bm{\theta}_{t-1})+\beta^{0}\nabla f_{B_{t}}(\bm{\theta}_{t})\}$
$=\beta^{t+1}\bm{m}_{-1}+(1-\beta)\sum_{i=0}^{t}\beta^{t-i}\nabla f_{B_{i}}(\bm{\theta}_{i})$
$=(1-\beta)\sum_{i=0}^{t}\beta^{t-i}\nabla f_{B_{i}}(\bm{\theta}_{i}),$
which, together with $\|\bm{\theta}\|^{2}=\langle\bm{\theta},\bm{\theta}\rangle$, implies that
$\left\|\boldsymbol{m}_{t}-(1-\beta)\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})\right\|^{2}$ $=(1-\beta)^{2}\left\|\sum_{i=0}^{t}\beta^{t-i}(\nabla f_{B_{i}}(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{i}))\right\|^{2}$
$=(1-\beta)^{2}\sum_{i=0}^{t}\sum_{j=0}^{t}\left\langle\beta^{t-i}(\nabla f_{B_{i}}(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{i})),\beta^{t-j}(\nabla f_{B_{j}}(\bm{\theta}_{j})-\nabla f(\bm{\theta}_{j}))\right\rangle.$
Let $i$ and $j$ satisfy $0\leq j<i\leq t$. Proposition A.1 and Assumptions (A2) and (A3) imply that
$\mathbb{E}\left[\left\langle\nabla f_{B_{i}}(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{i}),\nabla f_{B_{j}}(\bm{\theta}_{j})-\nabla f(\bm{\theta}_{j})\right\rangle\right]$
$=\mathbb{E}_{\bm{\xi}_{0}}\mathbb{E}_{\bm{\xi}_{1}}\cdots\mathbb{E}_{\bm{\xi}_{t}}\left[\langle\nabla f_{B_{i}}(\bm{\theta}_{i})-\mathbb{E}_{\bm{\xi}_{i}}\left[\nabla f_{B_{i}}(\bm{\theta}_{i})\right],\nabla f_{B_{j}}(\bm{\theta}_{j})-\mathbb{E}_{\bm{\xi}_{j}}[\nabla f_{B_{j}}(\bm{\theta}_{j})]\rangle\right]$
$=\mathbb{E}_{\bm{\xi}_{0}}\mathbb{E}_{\bm{\xi}_{1}}\cdots\mathbb{E}_{\bm{\xi}_{i}}\left[\langle\nabla f_{B_{i}}(\bm{\theta}_{i})-\mathbb{E}_{\bm{\xi}_{i}}[\nabla f_{B_{i}}(\bm{\theta}_{i})],\nabla f_{B_{j}}(\bm{\theta}_{j})-\mathbb{E}_{\bm{\xi}_{j}}[\nabla f_{B_{j}}(\bm{\theta}_{j})]\rangle\right]$
$=\mathbb{E}_{\bm{\xi}_{0}}\mathbb{E}_{\bm{\xi}_{1}}\cdots\mathbb{E}_{\bm{\xi}_{i-1}}\left[\langle\mathbb{E}_{\bm{\xi}_{i}}[\nabla f_{B_{i}}(\bm{\theta}_{i})]-\mathbb{E}_{\bm{\xi}_{i}}[\nabla f_{B_{i}}(\bm{\theta}_{i})],\nabla f_{B_{j}}(\bm{\theta}_{j})-\mathbb{E}_{\bm{\xi}_{j}}[\nabla f_{B_{j}}(\bm{\theta}_{j})]\rangle\right]$
$=0.$
A similar argument as in the case where $j<i$ ensures the above equation also holds for $i<j$. Hence, Proposition A.1 guarantees that, for all $t\in\mathbb{N}$,
$\mathbb{E}\left[\left\|\boldsymbol{m}_{t}-(1-\beta)\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})\right\|^{2}\right]$ $=(1-\beta)^{2}\sum_{i=0}^{t}\mathbb{E}\left[\left\langle\beta^{t-i}(\nabla f_{B_{i}}(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{i})),\beta^{t-i}(\nabla f_{B_{i}}(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{i}))\right\rangle\right]$
$=(1-\beta)^{2}\sum_{i=0}^{t}\beta^{2(t-i)}\mathbb{E}\left[\|\nabla f_{B_{i}}(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{i})\|^{2}\right]$
$\leq(1-\beta)^{2}\sum_{i=0}^{t}\beta^{2(t-i)}\frac{\sigma^{2}}{b_{i}},$
which completes the proof. ∎
A.2 Proofs of Theorems 3.3 and 3.4
Using Lemma A.2, we have the following.
Lemma A.3.
Suppose that Assumption 2.1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1. We define $(\bm{z}_{t})$ for all $t\in\{0\}\cup\mathbb{N}$ as
$\bm{z}_{t}:=\begin{cases}\bm{\theta}_{t}&\text{ }(t=0)\\ \frac{1}{1-\beta}\bm{\theta}_{t}-\frac{\beta}{1-\beta}\bm{\theta}_{t-1}&\text{}(t\geq 1).\end{cases}$
Then, for all $t\in\{0\}\cup\mathbb{N}$,
$\mathbb{E}[f(\boldsymbol{z}_{t+1})]$ $\leq\mathbb{E}[f(\boldsymbol{z}_{t})]+\eta\left[L\left\{\left(\frac{\beta}{1-\beta}\right)^{2}+\frac{3}{2}\right\}\eta-1\right]\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]+\frac{L\sigma^{2}\eta^{2}}{2}\left\{\beta^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+\frac{1}{b_{t}}\right\}$
$\quad+\left(\frac{1}{1-\beta}\right)^{2}L\eta^{2}{(1-\beta^{t})^{2}}\mathbb{E}\left[\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}\right],$
where $L:=\frac{1}{n}\sum_{i\in[n]}L_{i}$ is the Lipschitz constant of $\nabla f$ and we assume that $\sum_{i=0}^{-1}a_{i}:=0$ for some $a_{i}\in\mathbb{R}$.
Proof.
The descent lemma (Beck, 2017, Lemma 5.7) ensures that, for all $t\in\{0\}\cup\mathbb{N}$,
$\mathbb{E}_{\bm{\xi}_{t}}[f(\boldsymbol{z}_{t+1})]$ $\leq f(\boldsymbol{z}_{t})+\mathbb{E}_{\bm{\xi}_{t}}[\langle\nabla f(\boldsymbol{z}_{t}),\boldsymbol{z}_{t+1}-\boldsymbol{z}_{t}\rangle]+\frac{L}{2}\mathbb{E}_{\bm{\xi}_{t}}[\|\boldsymbol{z}_{t+1}-\boldsymbol{z}_{t}\|^{2}],$
which, together with $\boldsymbol{z}_{t+1}=\boldsymbol{z}_{t}-\eta\nabla f_{B_{t}}(\bm{\theta}_{t})$ (Liu et al., 2020, Lemma 3), implies that
$\mathbb{E}_{\bm{\xi}_{t}}[f(\boldsymbol{z}_{t+1})]$ $\leq f(\boldsymbol{z}_{t})-\eta\underbrace{\mathbb{E}_{\bm{\xi}_{t}}[\langle\nabla f(\boldsymbol{z}_{t}),\nabla f_{B_{t}}(\bm{\theta}_{t})\rangle]}_{X_{t}}+\frac{L\eta^{2}}{2}\underbrace{\mathbb{E}_{\bm{\xi}_{t}}[\|\nabla f_{B_{t}}(\bm{\theta}_{t})\|^{2}]}_{Y_{t}}.$ (7)
From Proposition A.1, we have that
$X_{t}=\langle\nabla f(\boldsymbol{z}_{t}),\mathbb{E}_{\bm{\xi}_{t}}[\nabla f_{B_{t}}(\bm{\theta}_{t})]\rangle=\langle\nabla f(\boldsymbol{z}_{t}),\nabla f(\bm{\theta}_{t})\rangle,$
which, together with the Cauchy–Schwarz inequality, Young’s inequality, and $L$-smoothness of $f$, implies that, for all $\rho>0$,
$-\eta X_{t}$ $=\langle\nabla f(\boldsymbol{z}_{t}),-\eta\nabla f(\bm{\theta}_{t})\rangle$
$=\langle\nabla f(\boldsymbol{z}_{t})-\nabla f(\bm{\theta}_{t}),-\eta\nabla f(\bm{\theta}_{t})\rangle-\eta\|\nabla f(\bm{\theta}_{t})\|^{2}$
$\leq(\sqrt{\eta}\|\nabla f(\boldsymbol{z}_{t})-\nabla f(\bm{\theta}_{t})\|)(\sqrt{\eta}\|\nabla f(\bm{\theta}_{t})\|)-\eta\|\nabla f(\bm{\theta}_{t})\|^{2}$
$\leq\frac{\rho\eta}{2}\|\nabla f(\boldsymbol{z}_{t})-\nabla f(\boldsymbol{\theta}_{t})\|^{2}+\frac{\eta}{2\rho}\|\nabla f(\bm{\theta}_{t})\|^{2}-\eta\|\nabla f(\bm{\theta}_{t})\|^{2}$
$\leq\frac{\rho\eta L^{2}}{2}\|\boldsymbol{z}_{t}-\boldsymbol{\theta}_{t}\|^{2}+\eta\left(\frac{1}{2\rho}-1\right)\|\nabla f(\bm{\theta}_{t})\|^{2}.$
The definitions of $\bm{z}_{t}$ and $\bm{\theta}_{t}$ ($=\bm{\theta}_{t-1}-\eta\bm{m}_{t-1}$) ensure that, for all $t\geq 1$,
$\boldsymbol{z}_{t}=\frac{1}{1-\beta}\boldsymbol{\theta}_{t}-\frac{\beta}{1-\beta}(\boldsymbol{\theta}_{t}+\eta\boldsymbol{m}_{t-1})=\boldsymbol{\theta}_{t}-\frac{\beta}{1-\beta}\eta\boldsymbol{m}_{t-1}.$
From $\bm{m}_{-1}:=\bm{0}$ and $\bm{z}_{0}=\bm{\theta}_{0}$, we have that, for all $t\in\{0\}\cup\mathbb{N}$,
$\boldsymbol{z}_{t}=\boldsymbol{\theta}_{t}-\frac{\beta}{1-\beta}\eta\boldsymbol{m}_{t-1}.$
Accordingly, we have that
$-\eta X_{t}\leq\frac{\rho\eta^{3}L^{2}}{2}\left(\frac{\beta}{1-\beta}\right)^{2}\|\boldsymbol{m}_{t-1}\|^{2}+\eta\left(\frac{1}{2\rho}-1\right)\|\nabla f(\bm{\theta}_{t})\|^{2}.$ (8)
From $\|\bm{\theta}_{1}+\bm{\theta}_{2}\|^{2}\leq 2\|\bm{\theta}_{1}\|^{2}+2\|\bm{\theta}_{2}\|^{2}$, we have that
$\|\boldsymbol{m}_{t-1}\rVert^{2}$ $\leq 2\left\|\boldsymbol{m}_{t-1}-(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}+2\underbrace{\left\|(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}}_{Z_{t}}.$ (9)
Moreover, for all $t\geq 2$,
$\frac{1}{(1-\beta^{t-1})^{2}}Z_{t}$ $\leq 2\|\nabla f(\bm{\theta}_{t})\|^{2}+2\underbrace{\left\|\frac{1-\beta}{1-\beta^{t-1}}\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}}_{W_{t}}.$ (10)
Meanwhile, we also have that
$\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}$
$=\left\|\frac{1-\beta}{1-\beta^{t}}\left(\beta^{t}\nabla f(\bm{\theta}_{0})+\beta^{t-1}\nabla f(\bm{\theta}_{1})+\cdots+\beta^{t-(t-1)}\nabla f(\bm{\theta}_{t-1})+\nabla f(\bm{\theta}_{t})\right)-\nabla f(\bm{\theta}_{t})\right\rVert^{2}$
$=\left\|\frac{1-\beta}{1-\beta^{t}}\left(\beta^{t}\nabla f(\bm{\theta}_{0})+\beta^{t-1}\nabla f(\bm{\theta}_{1})+\cdots+\beta^{t-(t-1)}\nabla f(\bm{\theta}_{t-1})\right)+\left(\frac{1-\beta}{1-\beta^{t}}-1\right)\nabla f(\bm{\theta}_{t})\right\rVert^{2}$
$=\left\|\frac{1-\beta}{1-\beta^{t}}\left(\beta^{t}\nabla f(\bm{\theta}_{0})+\beta^{t-1}\nabla f(\bm{\theta}_{1})+\cdots+\beta^{t-(t-1)}\nabla f(\bm{\theta}_{t-1})\right)-\frac{\beta-\beta^{t}}{1-\beta^{t}}\nabla f(\bm{\theta}_{t})\right\rVert^{2}$
$=\left\|\frac{1-\beta}{1-\beta^{t}}\beta\left(\beta^{t-1}\nabla f(\bm{\theta}_{0})+\beta^{t-2}\nabla f(\bm{\theta}_{1})+\cdots+\beta^{t-(t-1)}\nabla f(\bm{\theta}_{t-2})+\nabla f(\bm{\theta}_{t-1})\right)-\frac{1-\beta^{t-1}}{1-\beta^{t}}\beta\nabla f(\bm{\theta}_{t})\right\rVert^{2}$
$=\beta^{2}\left(\frac{1-\beta^{t-1}}{1-\beta^{t}}\right)^{2}\left\|\frac{1-\beta}{1-\beta^{t-1}}\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}$
$=\beta^{2}\left(\frac{1-\beta^{t-1}}{1-\beta^{t}}\right)^{2}W_{t},$
which implies that, for all $t\geq 2$,
$W_{t}=\frac{(1-\beta^{t})^{2}}{\beta^{2}(1-\beta^{t-1})^{2}}\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}.$ (11)
From (9), (10), and (11),
$\|\boldsymbol{m}_{t-1}\rVert^{2}$ $\leq 2\left\|\boldsymbol{m}_{t-1}-(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}+2\Bigg{\{}2(1-\beta^{t-1})^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}$
$\quad+2(1-\beta^{t-1})^{2}\frac{(1-\beta^{t})^{2}}{\beta^{2}(1-\beta^{t-1})^{2}}\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}\Bigg{\}}$
$=2\left\|\boldsymbol{m}_{t-1}-(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}+4(1-\beta^{t-1})^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}$
$\quad+\frac{4(1-\beta^{t})^{2}}{\beta^{2}}\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}$
Hence, (8) ensures that
$-\eta X_{t}$ $\leq\frac{\rho\eta^{3}L^{2}}{2}\left(\frac{\beta}{1-\beta}\right)^{2}\Bigg{\{}2\left\|\boldsymbol{m}_{t-1}-(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}+4(1-\beta^{t-1})^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}$
$\quad+\frac{4(1-\beta^{t})^{2}}{\beta^{2}}\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}\Bigg{\}}+\eta\left(\frac{1}{2\rho}-1\right)\|\nabla f(\bm{\theta}_{t})\|^{2}$
$=\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}\left\|\boldsymbol{m}_{t-1}-(1-\beta)\sum_{i=0}^{t-1}\beta^{t-1-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}+2\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}(1-\beta^{t-1})^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}$
$\quad+2\rho\eta^{3}L^{2}\left(\frac{1}{1-\beta}\right)^{2}(1-\beta^{t})^{2}\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}+\eta\left(\frac{1}{2\rho}-1\right)\|\nabla f(\bm{\theta}_{t})\|^{2}.$
Let us take the total expectation on both sides of the above inequality. Lemma A.2 then guarantees that, for all $t\geq 2$ and for all $\rho>0$,
$\begin{split}-\eta\mathbb{E}[X_{t}]&\leq\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+2\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}(1-\beta^{t-1})^{2}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\\ &\quad+2\rho\eta^{3}L^{2}\left(\frac{1}{1-\beta}\right)^{2}(1-\beta^{t})^{2}\mathbb{E}\left[\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}\right]+\eta\left(\frac{1}{2\rho}-1\right)\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\\ &\leq\rho\eta^{3}L^{2}\beta^{2}\sigma^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+2\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\\ &\quad+2\rho\eta^{3}L^{2}\left(\frac{1}{1-\beta}\right)^{2}{(1-\beta^{t})^{2}}\mathbb{E}\left[\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}\right]+\eta\left(\frac{1}{2\rho}-1\right)\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}].\end{split}$ (12)
Moreover, Proposition A.1 guarantees that
$\begin{split}\mathbb{E}[Y_{t}]&=\mathbb{E}_{\bm{\xi}_{t}}\left[\left\|\nabla f_{B_{t}}(\bm{\theta}_{t})-\nabla f(\bm{\theta}_{t})+\nabla f(\bm{\theta}_{t})\right\|^{2}\Big{|}\hat{\bm{\xi}}_{t-1}\right]\\ &=\mathbb{E}[\|\nabla f_{B_{t}}(\bm{\theta}_{t})-\nabla f(\bm{\theta}_{t})\|^{2}]+2\mathbb{E}[\langle\nabla f_{B_{t}}(\bm{\theta}_{t})-\nabla f(\bm{\theta}_{t}),\nabla f(\bm{\theta}_{t})\rangle]+\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\\ &\leq\frac{\sigma^{2}}{b_{t}}+\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}].\end{split}$ (13)
Therefore, from (7), (12), and (13), for all $t\geq 2$ and for all $\rho>0$,
$\mathbb{E}[f(\boldsymbol{z}_{t+1})]$ $\leq\mathbb{E}[f(\boldsymbol{z}_{t})]-\eta\mathbb{E}[X_{t}]+\frac{L\eta^{2}}{2}\mathbb{E}[Y_{t}]$
$\leq\mathbb{E}[f(\boldsymbol{z}_{t})]+\rho\eta^{3}L^{2}\beta^{2}\sigma^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+2\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}+\eta\left(\frac{1}{2\rho}-1\right)\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
$\quad+2\rho\eta^{3}L^{2}\left(\frac{1}{1-\beta}\right)^{2}{(1-\beta^{t})^{2}}\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}+\frac{L\eta^{2}}{2}\left(\frac{\sigma^{2}}{b_{t}}+\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\right)$
$=\mathbb{E}[f(\boldsymbol{z}_{t})]+\underbrace{\left\{2\rho\eta^{3}L^{2}\left(\frac{\beta}{1-\beta}\right)^{2}+\eta\left(\frac{1}{2\rho}-1\right)+\frac{L\eta^{2}}{2}\right\}}_{A}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
$\quad+L\eta^{2}\sigma^{2}\underbrace{\left(\rho\eta L\beta^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+\frac{1}{2b_{t}}\right)}_{B_{t}}$
$\quad+\underbrace{2\rho\eta^{3}L^{2}}_{C}\left(\frac{1}{1-\beta}\right)^{2}{(1-\beta^{t})^{2}}\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}.$
The setting $\rho:=\frac{1}{2L\eta}$ implies that, for all $t\geq 2$,
$A$ $=\frac{L^{2}\eta^{3}}{L\eta}\left(\frac{\beta}{1-\beta}\right)^{2}+\eta\left(L\eta-1\right)+\frac{L\eta^{2}}{2}=L\eta^{2}\left(\frac{\beta}{1-\beta}\right)^{2}+\eta\left(L\eta-1\right)+\frac{L\eta^{2}}{2}$
$=L\left\{\left(\frac{\beta}{1-\beta}\right)^{2}+\frac{3}{2}\right\}\eta^{2}-\eta,$
$B_{t}$ $=\frac{L\eta\beta^{2}}{2L\eta}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+\frac{1}{2b_{t}}=\frac{1}{2}\left(\beta^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+\frac{1}{b_{t}}\right),\text{ }C=\frac{2L^{2}\eta^{3}}{2L\eta}=L\eta^{2}.$
When $t=0$, from $\|\bm{m}_{-1}\|=0$ and $\sum_{i=0}^{-1}a_{i}:=0$, the assertion in Lemma A.3 holds. When $t=1$, assuming $\frac{1}{1-\beta^{0}}:=1$, the assertion in Lemma A.3 again holds. This completes the proof. ∎
Using Lemma A.3, we have the following lemma.
Lemma A.4.
Suppose that Assumption 2.1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1 with $\eta>0$ satisfying
$\eta\leq\frac{1-\beta}{2\sqrt{2}\sqrt{\beta+\beta^{2}}L}.$
Let $(\bm{z}_{t})$ be the sequence defined as in Lemma A.3 and define $L_{t}\in\mathbb{R}$ for all $t\in\{0\}\cup\mathbb{N}$ as
$L_{t}:=\begin{cases}f(\boldsymbol{z}_{0})-f^{\star}&\text{ }(t=0)\\ f(\boldsymbol{z}_{1})-f^{\star}&\text{ }(t=1)\\ f(\boldsymbol{z}_{t})-f^{\star}+\sum_{i=1}^{t-1}c_{i}\left\|\boldsymbol{\theta}_{t+1-i}-\boldsymbol{\theta}_{t-i}\right\rVert^{2}&\text{ }(t\geq 2),\end{cases}$
where $f^{\star}$ is the minimum value of $f$ over $\mathbb{R}^{d}$ and $(c_{i})\subset\mathbb{R}_{++}$ is defined by
$c_{1}=\dfrac{(\beta+\beta^{2})L^{3}\eta^{2}}{(1-\beta)^{2}\{(1-\beta)^{2}-4(\beta+\beta^{2})L^{2}\eta^{2}\}}\text{ and }c_{i+1}=c_{i}-\left(4c_{1}\eta^{2}+\frac{L\eta^{2}}{{(1-\beta)^{2}}}\right)\beta^{i}\left(i+\frac{\beta}{1-\beta}\right)L^{2}\quad(i\in[t-1]).$
Then, for all $t\in\{0\}\cup\mathbb{N}$,
$\mathbb{E}[L_{t+1}-L_{t}]$ $\leq\eta\left[L\left\{\left(\frac{\beta}{1-\beta}\right)^{2}+\frac{3}{2}\right\}\eta-1+4c_{1}\eta\right]\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
$\quad+\frac{L\sigma^{2}\eta^{2}}{2}\left\{\beta^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+\frac{1}{b_{t}}\right\}+2c_{1}\eta^{2}(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}},$
where we assume that $\sum_{i=0}^{-1}a_{i}:=0$ for some $a_{i}\in\mathbb{R}$.
Proof.
Let $t\geq 2$. The definition of $L_{t}$ implies that
$\mathbb{E}[L_{t+1}-L_{t}]$ $=\mathbb{E}[f(\boldsymbol{z}_{t+1})-f(\boldsymbol{z}_{t})]+\mathbb{E}\left[\sum_{i=1}^{t}c_{i}\left\|\boldsymbol{\theta}_{t+2-i}-\boldsymbol{\theta}_{t+1-i}\right\rVert^{2}-\sum_{i=1}^{t-1}c_{i}\left\|\boldsymbol{\theta}_{t+1-i}-\boldsymbol{\theta}_{t-i}\right\rVert^{2}\right]$
$=\mathbb{E}[f(\boldsymbol{z}_{t+1})-f(\boldsymbol{z}_{t})]+\mathbb{E}[c_{1}\left\|\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}_{t}\right\rVert^{2}]+\mathbb{E}\left[\sum_{i=1}^{t-1}c_{i+1}\left\|\boldsymbol{\theta}_{t+1-i}-\boldsymbol{\theta}_{t-i}\right\rVert^{2}-\sum_{i=1}^{t-1}c_{i}\left\|\boldsymbol{\theta}_{t+1-i}-\boldsymbol{\theta}_{t-i}\right\rVert^{2}\right]$
$=\mathbb{E}[f(\boldsymbol{z}_{t+1})-f(\boldsymbol{z}_{t})]+\mathbb{E}[c_{1}\left\|\boldsymbol{\theta}_{t+1}-\boldsymbol{\theta}_{t}\right\rVert^{2}]+\sum_{i=1}^{t-1}(c_{i+1}-c_{i})\mathbb{E}[\left\|\boldsymbol{\theta}_{t+1-i}-\boldsymbol{\theta}_{t-i}\right\rVert^{2}].$
Lemma A.3 thus ensures that
$\mathbb{E}[L_{t+1}-L_{t}]$ $\leq\eta\left[L\left\{\left(\frac{\beta}{1-\beta}\right)^{2}+\frac{3}{2}\right\}\eta-1\right]\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]+\frac{L\sigma^{2}\eta^{2}}{2}\left\{\beta^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+\frac{1}{b_{t}}\right\}$
$\quad+\left(\frac{1}{1-\beta}\right)^{2}L\eta^{2}{(1-\beta^{t})^{2}}\mathbb{E}\left[\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}\right]$
$\quad+c_{1}\eta^{2}\mathbb{E}[\left\|\boldsymbol{m}_{t}\right\rVert^{2}]+\sum_{i=1}^{t-1}(c_{i+1}-c_{i})\mathbb{E}[\left\|\boldsymbol{\theta}_{t+1-i}-\boldsymbol{\theta}_{t-i}\right\rVert^{2}].$ (14)
A similar discussion to the one for showing (9) and (10) ensures that
$\|\boldsymbol{m}_{t}\rVert^{2}$ $\leq 2\left\|\boldsymbol{m}_{t}-(1-\beta)\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})\right\rVert^{2}$
$\quad+2\left\{2(1-\beta^{t})^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}+2(1-\beta^{t})^{2}\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}\right\},$
which, together with Lemma A.2 and $1-\beta^{t}\leq 1$, implies that
$\mathbb{E}[\|\bm{m}_{t}\|^{2}]$ $\leq 2(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}+4(1-\beta^{t})^{2}\|\nabla f(\bm{\theta}_{t})\|^{2}+4(1-\beta^{t})^{2}\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}$
$\leq 2(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}+4\|\nabla f(\bm{\theta}_{t})\|^{2}+4(1-\beta^{t})^{2}\underbrace{\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}}_{V_{t}}.$
Lemma 2 in (Liu et al., 2020) guarantees that
$\mathbb{E}[V_{t}]\leq\sum_{i=1}^{t-1}a_{t,i}\mathbb{E}[\|\bm{\theta}_{i+1}-\bm{\theta}_{i}\|^{2}],\text{ where }a_{t,i}:=\frac{L^{2}\beta^{t-i}}{1-\beta^{t}}\left(t-i+\frac{\beta}{1-\beta}\right),$
which implies that
$\mathbb{E}[V_{t}]\leq\sum_{i=1}^{t-1}a_{t,t-i}\mathbb{E}[\|\bm{\theta}_{t+1-i}-\bm{\theta}_{t-i}\|^{2}],\text{ where }a_{t,t-i}:=\frac{L^{2}\beta^{i}}{1-\beta^{t}}\left(i+\frac{\beta}{1-\beta}\right).$ (15)
Moreover, $c_{1}>0$ when $\eta\leq\frac{1-\beta}{2\sqrt{2}L\sqrt{\beta+\beta^{2}}}$. Hence, (A.2) ensures that
$\mathbb{E}[L_{t+1}-L_{t}]$ $\leq\eta\left[L\left\{\left(\frac{\beta}{1-\beta}\right)^{2}+\frac{3}{2}\right\}\eta-1\right]\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
$\quad+\frac{L\sigma^{2}\eta^{2}}{2}\left\{\beta^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+\frac{1}{b_{t}}\right\}+\left(\frac{1}{1-\beta}\right)^{2}L\eta^{2}{(1-\beta^{t})^{2}}\mathbb{E}[V_{t}]$
$\quad+c_{1}\eta^{2}\left\{2(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}+4\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]+4(1-\beta^{t})^{2}\mathbb{E}[V_{t}]\right\}$
$\quad+\sum_{i=1}^{t-1}(c_{i+1}-c_{i})\mathbb{E}[\left\|\boldsymbol{\theta}_{t+1-i}-\boldsymbol{\theta}_{t-i}\right\rVert^{2}]$
$=\eta\left[L\left\{\left(\frac{\beta}{1-\beta}\right)^{2}+\frac{3}{2}\right\}\eta-1+4c_{1}\eta\right]\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
$\quad+\frac{L\sigma^{2}\eta^{2}}{2}\left\{\beta^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+\frac{1}{b_{t}}\right\}+2c_{1}\eta^{2}(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}$
$\quad+\sum_{i=1}^{t-1}\underbrace{\left\{\left(\frac{1}{1-\beta}\right)^{2}L\eta^{2}{(1-\beta^{t})^{2}}a_{t,t-i}+4c_{1}\eta^{2}(1-\beta^{t})^{2}a_{t,t-i}+(c_{i+1}-c_{i})\right\}}_{N_{t,i}}\mathbb{E}[\left\|\boldsymbol{\theta}_{t+1-i}-\boldsymbol{\theta}_{t-i}\right\rVert^{2}].$
Finally, we prove that $N_{t,i}\leq 0$. From the definition of $a_{t,t-i}$ in (15) and
${c_{i+1}=c_{i}-\left(4c_{1}\eta^{2}+\frac{L\eta^{2}}{(1-\beta)^{2}}\right)\beta^{i}\left(i+\frac{\beta}{{1-\beta}}\right)L^{2},}$
we have that
$N_{t,i}$ $=\left(\frac{1}{1-\beta}\right)^{2}L\eta^{2}{(1-\beta^{t})^{2}}a_{t,t-i}+4c_{1}\eta^{2}(1-\beta^{t})^{2}a_{t,t-i}+(c_{i+1}-c_{i})$
=$=\left\{\left(\frac{1}{1-\beta}\right)^{2}L\eta^{2}{(1-\beta^{t})^{2}}+4c_{1}\eta^{2}(1-\beta^{t})^{2}\right\}\frac{L^{2}\beta^{i}}{1-\beta^{t}}\left(i+\frac{\beta}{1-\beta}\right)-\left(4c_{1}\eta^{2}+\frac{L\eta^{2}}{{(1-\beta)^{2}}}\right)\beta^{i}\left(i+\frac{\beta}{1-\beta}\right)L^{2}$
$=L^{2}\eta^{2}\beta^{i}\left(i+\frac{\beta}{1-\beta}\right)\left[\left\{\frac{1-\beta^{t}}{(1-\beta)^{2}}L+4c_{1}(1-\beta^{t})\right\}-\left\{4c_{1}+\frac{L}{(1-\beta)^{2}}\right\}\right]$
$=L^{2}\eta^{2}\beta^{i}\left(i+\frac{\beta}{1-\beta}\right)\left[\frac{L}{(1-\beta)^{2}}(1-\beta^{t}-1)+4c_{1}(1-\beta^{t}-1)\right]$
$=-L^{2}\eta^{2}\beta^{i}\left(i+\frac{\beta}{1-\beta}\right)\left[\frac{L}{(1-\beta)^{2}}+4c_{1}\right]\beta^{t}.$
From
$\eta\leq\frac{1-\beta}{2\sqrt{2}L\sqrt{\beta+\beta^{2}}}\text{ and }c_{1}=\dfrac{(\beta+\beta^{2})L^{3}\eta^{2}}{(1-\beta)^{2}\{(1-\beta)^{2}-4(\beta+\beta^{2})L^{2}\eta^{2}\}},$ (16)
we have that
$c_{1}=\frac{\eta^{2}L^{3}\frac{\beta+\beta^{2}}{(1-\beta)^{4}}}{1-4\eta^{2}L^{2}\frac{\beta+\beta^{2}}{(1-\beta)^{2}}}>0.$
Accordingly,
$N_{t,i}=-L^{2}\eta^{2}\beta^{i}\left(i+\frac{\beta}{1-\beta}\right)\left[\frac{L}{(1-\beta)^{2}}+4c_{1}\right]\beta^{t}<0.$
This completes the proof. ∎
Lemma A.4 leads to the following.
Lemma A.5.
Suppose that Assumption 2.1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1 with $\eta>0$ satisfying
$\eta\leq\max\left\{\frac{1-\beta}{2\sqrt{2}\sqrt{\beta+\beta^{2}}L},\frac{(1-\beta)^{2}}{(5\beta^{2}-6\beta+5)L}\right\}.$
Then, for all $T\geq 1$,
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\leq\frac{2(f(\bm{\theta}_{0})-f^{\star})}{\eta T}+\frac{2L\eta\sigma^{2}}{T}\sum_{t=0}^{T-1}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}.$
Proof.
Lemma A.4 guarantees that, for all $t\in\{0\}\cup\mathbb{N}$,
$\underbrace{-\eta\left[L\left\{\left(\frac{\beta}{1-\beta}\right)^{2}+\frac{3}{2}\right\}\eta-1+4c_{1}\eta\right]}_{D}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
$\quad\leq\mathbb{E}[L_{t+1}-L_{t}]+\underbrace{\frac{L\sigma^{2}\eta^{2}}{2}\left\{\beta^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+\frac{1}{b_{t}}\right\}+2c_{1}\eta^{2}(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}}_{U_{t}},$
where $\beta\in[0,1)$, and $\eta$ and $c_{1}$ satisfy (16). Summing the above inequality from $t=0$ to $t=T-1$ gives that
$D\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\leq\sum_{t=0}^{T-1}\mathbb{E}[L_{t}-L_{t+1}]+\sum_{t=0}^{T-1}U_{t}=\mathbb{E}[L_{0}-L_{T}]+\sum_{t=0}^{T-1}U_{t},$
which, together with $L_{T}\geq 0$, implies that
$D\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\leq L_{0}+\sum_{t=0}^{T-1}U_{t}.$ (17)
From (16), we have that
$c_{1}=\frac{\eta^{2}L^{3}\frac{\beta+\beta^{2}}{(1-\beta)^{4}}}{1-4\eta^{2}L^{2}\frac{\beta+\beta^{2}}{(1-\beta)^{2}}}\text{ and }\eta^{2}L^{2}\frac{\beta+\beta^{2}}{(1-\beta)^{2}}\leq\frac{1}{8},$
which implies that
$c_{1}\leq\frac{L}{8(1-\beta)^{2}}\left(1-\frac{4}{8}\right)^{-1}=\frac{L}{4(1-\beta)^{2}}.$ (18)
Accordingly, from (18) and $\eta\leq\frac{(1-\beta)^{2}}{L(5\beta^{2}-6\beta+5)}$, we have that
$\begin{split}D&=-L\left\{\left(\frac{\beta}{1-\beta}\right)^{2}+\frac{3}{2}\right\}\eta^{2}+\eta-4c_{1}\eta^{2}\geq-L\left\{\left(\frac{\beta}{1-\beta}\right)^{2}+\frac{3}{2}\right\}\eta^{2}+\eta-\frac{L\eta^{2}}{(1-\beta)^{2}}\\ &=-L\eta^{2}\frac{5\beta^{2}-6\beta+5}{2(1-\beta)^{2}}+\eta\geq-\frac{\eta}{2}+\eta=\frac{\eta}{2}>0.\end{split}$ (19)
Moreover, from (18), we have that
$\begin{split}U_{t}&=\frac{L\sigma^{2}\eta^{2}}{2}\left\{\beta^{2}\sum_{i=0}^{t-1}\frac{\beta^{2(t-1-i)}}{b_{i}}+\frac{1}{b_{t}}\right\}+2c_{1}\eta^{2}(1-\beta)^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}\\ &=\sigma^{2}\left\{\frac{L\eta^{2}}{2}+2c_{1}\eta^{2}(1-\beta)^{2}\right\}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}\\ &\leq\sigma^{2}\left(\frac{L\eta^{2}}{2}+\frac{L\eta^{2}}{2}\right)\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}=L\eta^{2}\sigma^{2}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}.\end{split}$ (20)
Therefore, (17), (19), and (20) ensure that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\leq\frac{L_{0}}{DT}+\frac{1}{T}\sum_{t=0}^{T-1}U_{t}\leq\frac{2L_{0}}{\eta T}+\frac{2L\eta\sigma^{2}}{T}\sum_{t=0}^{T-1}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}.$
This completes the proof. ∎
Proof of Theorem 3.3.
Let $b_{t}$ be defined by (4), i.e., for all $m\in[0:M]$ and all $t\in S_{m}=\mathbb{N}\cap[\sum_{k=0}^{m-1}K_{k}E_{k},\sum_{k=0}^{m}K_{k}E_{k})$ ($S_{0}:=\mathbb{N}\cap[0,K_{0}E_{0})$),
$b_{t}=\delta^{m\left\lceil\frac{t}{\sum_{k=0}^{m}K_{k}E_{k}}\right\rceil}b_{0},$
which implies that $b_{j}=\delta^{j}b_{0}$ ($j\in S_{j}$) and
$(b_{0},b_{1},\cdots,b_{M})=(\underbrace{b_{0},b_{0},\cdots,b_{0}}_{K_{0}E_{0}},\underbrace{\delta b_{0},\delta b_{0},\cdots,\delta b_{0}}_{K_{1}E_{1}},\cdots,\underbrace{\delta^{M}b_{0},\delta^{M}b_{0},\cdots,\delta^{M}b_{0}}_{K_{M}E_{M}}),$
where $T=\sum_{m=0}^{M}K_{m}E_{m}$. Define $K_{\max}:=\max\{K_{m}\colon m\in[0:M]\}$ and $E_{\max}:=\max\{E_{m}\colon m\in[0:M]\}$. Then, we have that
$\sum_{t=0}^{T-1}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}$ $=\sum_{t=0}^{T-1}\sum_{i=0}^{t}\beta^{2(t-i)}\frac{K_{i}E_{i}}{\delta^{i}b_{0}}\leq\frac{K_{\max}E_{\max}}{b_{0}}\sum_{t=0}^{T-1}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{\delta^{i}}$
$=\frac{K_{\max}E_{\max}}{b_{0}}\sum_{t=0}^{T-1}\beta^{2t}\sum_{i=0}^{t}\frac{1}{(\beta^{2}\delta)^{i}}=\frac{K_{\max}E_{\max}}{b_{0}}\sum_{t=0}^{T-1}\beta^{2t}\frac{1-(\frac{1}{\beta^{2}\delta})^{t+1}}{1-\frac{1}{\beta^{2}\delta}}$
$=\frac{K_{\max}E_{\max}}{b_{0}}\sum_{t=0}^{T-1}\beta^{2t}\frac{1-(\frac{1}{\beta^{2}\delta})^{t+1}}{1-\frac{1}{\beta^{2}\delta}}=\frac{K_{\max}E_{\max}\delta}{b_{0}(\beta^{2}\delta-1)}\sum_{t=0}^{T-1}\left\{\beta^{2(t+1)}-\frac{1}{\delta^{t+1}}\right\},$
which implies that
$\sum_{t=0}^{T-1}\sum_{i=0}^{t}\frac{\beta^{2(t-i)}}{b_{i}}\leq\frac{K_{\max}E_{\max}\delta}{b_{0}(\beta^{2}\delta-1)}\left\{\frac{\beta^{2}(1-\beta^{2T})}{1-\beta^{2}}-\frac{1-(\frac{1}{\delta})^{T}}{\delta-1}\right\}\leq\frac{K_{\max}E_{\max}\delta}{b_{0}(\beta^{2}\delta-1)}\left(\frac{\beta^{2}}{1-\beta^{2}}-\frac{1}{\delta-1}\right).$
This completes the proof. ∎
Proof of Theorem 3.4.
NSHB with $\eta=\frac{\alpha}{1-\beta}$ coincides with SHB defined by (2) (Section 2.2). Hence, Theorem 3.3 leads to Theorem 3.4. ∎
A.3 Proofs of Theorems 3.1 and 3.2
Using Lemma A.2 and the proof of Lemma A.3 with $\rho=\frac{1-\beta}{2L\eta}$, we have the following lemma. Hence, we omit the proof of Lemma A.6.
Lemma A.6.
Suppose that Assumption 2.1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1. Let $(\bm{z}_{t})$ be the sequence defined as in Lemma A.3. Then, for all $t\in\{0\}\cup\mathbb{N}$,
$\mathbb{E}[f(\boldsymbol{z}_{t+1})]$ $\leq\mathbb{E}[f(\boldsymbol{z}_{t})]+\eta\left\{L\left(\frac{1+\beta^{2}}{1-\beta}+\frac{1}{2}\right)\eta-1\right\}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]$
$\quad+\frac{L\sigma^{2}\eta^{2}}{2}\left(\frac{\beta^{2}}{1+\beta}+1\right)+\frac{(1-\beta^{t})^{2}}{1-\beta}L\eta^{2}\mathbb{E}\left[\left\|\frac{1-\beta}{1-\beta^{t}}\sum_{i=0}^{t}\beta^{t-i}\nabla f(\bm{\theta}_{i})-\nabla f(\bm{\theta}_{t})\right\rVert^{2}\right],$
where $L:=\frac{1}{n}\sum_{i\in[n]}L_{i}$ is the Lipschitz constant of $\nabla f$ and we assume that $\sum_{i=0}^{-1}a_{i}:=0$ for some $a_{i}\in\mathbb{R}$.
Using Lemma A.6 and the proof of Lemma A.4 with $\rho=\frac{1-\beta}{2L\eta}$, we have the following lemma. Hence, we omit the proof of Lemma A.7.
Lemma A.7.
Suppose that Assumption 2.1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1 with $\eta>0$ satisfying
$\eta\leq\frac{1-\beta}{2\sqrt{2}\sqrt{\beta+\beta^{2}}L}.$
Let $(\bm{z}_{t})$ be the sequence defined as in Lemma A.3 and let $L_{t}\in\mathbb{R}$ be defined as in Lemma A.4, where $(c_{i})\subset\mathbb{R}_{++}$ is defined by
$c_{1}=\dfrac{(\beta+\beta^{2})L^{3}\eta^{2}}{(1-\beta)\{(1-\beta)^{2}-4(\beta+\beta^{2})L^{2}\eta^{2}\}}\text{ and }c_{i+1}=c_{i}-\left(4c_{1}\eta^{2}+\frac{L\eta^{2}}{1-\beta}\right)\beta^{i}\left(i+\frac{\beta}{1-\beta}\right)L^{2}\quad(i\in[t-1]).$
Then, for all $t\in\{0\}\cup\mathbb{N}$,
$\mathbb{E}[L_{t+1}-L_{t}]$ $\leq\eta\left\{\frac{(3-\beta+\beta^{2})L\eta}{2(1-\beta)}-1+4c_{1}\eta\right\}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]+\left\{\left(\frac{\beta^{2}}{1+\beta}+1\right)\frac{L}{2}+\frac{2c_{1}(1-\beta)}{1+\beta}\right\}\frac{\eta^{2}\sigma^{2}}{b}.$
Lemma A.7 and the proof of Lemma A.5 with $\rho=\frac{1-\beta}{2L\eta}$ lead to the following.
Lemma A.8.
Suppose that Assumption 2.1 holds and $(\bm{\theta}_{t})$ is the sequence generated by Algorithm 1 with $\eta>0$ satisfying
$\eta\leq\max\left\{\frac{1-\beta}{2\sqrt{2}\sqrt{\beta+\beta^{2}}L},\frac{1-\beta}{(5-\beta+2\beta^{2})L}\right\}.$
Then, for all $T\geq 1$,
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bm{\theta}_{t})\|^{2}]\leq\frac{2(f(\bm{\theta}_{0})-f^{\star})}{\eta T}+\frac{L\eta\sigma^{2}}{b}\left\{\frac{\beta+3\beta^{2}}{2(1+\beta)}+1\right\}.$
Proofs of Theorems 3.1 and 3.2.
Lemma A.8 leads to the assertion in Theorem 3.1. NSHB with $\eta=\frac{\alpha}{1-\beta}$ coincides with SHB defined by (2) (Section 2.2). Hence, Theorem 3.1 leads to Theorem 3.2. ∎
A.4 Training ResNet-18 on Tiny ImageNet
(a) Learning rate and batch size versus epochs
(b) Full gradient norm versus epochs
(c) Empirical loss versus epochs
(d) Test accuracy score versus epochs
Figure 4: (a) Schedulers for each optimizer with constant learning rates and constant batch size, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on Tiny ImageNet dataset.
(a) Learning rate and batch size versus epochs
(b) Full gradient norm versus epochs
(c) Empirical loss versus epochs
(d) Test accuracy score versus epochs
Figure 5: (a) Schedulers for each optimizer with constant learning rates and doubly increasing batch size every 25 epochs, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on Tiny ImageNet dataset.
(a) Learning rate and batch size versus epochs
(b) Full gradient norm versus epochs
(c) Empirical loss versus epochs
(d) Test accuracy score versus epochs
Figure 6: (a) Schedulers for each optimizer with constant learning rates and quadrupling increasing batch size every 50 epochs, (b) Full gradient norm of empirical loss for training, (c) Empirical loss value for training, and (d) Accuracy score for test to train ResNet-18 on Tiny ImageNet dataset.